New lithographic equipment, acquired at a cost of $905,600 on March 1 at the beginning of a fiscal year, has an estimated useful life of five years and an estimated residual value of $101,880. The manager requested information regarding the effect of alternative methods on the amount of depreciation expense each year. On the basis of the data presented to the manager, the double-declining-balance method was selected. In the first week of the fifth year, on March 4, the equipment was sold for $149,286. Required: 1. Determine the annual depreciation expense for each of the estimated five years of use, the accumulated depreciation at the end of each year, and the book value of the equipment at the end of each year by (a) the straight-line method and (b) the double-declining-balance method. Round your answers to the nearest whole dollar. 2. Journalize the entry to record the sale assuming the manager chose the double-declining-balance method. Refer to the Chart of Accounts for exact wording of account titles. 3. Journalize the entry to record the sale in (2), assuming that the equipment was sold for $103,386 instead of $149,286. Refer to the Chart of Accounts for exact wording of account titles. Determine the annual depreciation expense for each of the estimated five years of use, the accumulated depreciation at the end of each year, and the book value of the equipment at the end of each year by (a) the straight-line method and (b) the double-declining-balance method. Round your answers to the nearest whole dollar. Enter all amounts as a positive numbers. a. Straight-line method Accumulated Depreciation, Year Depreciation Expense End of Year Book Value, End of Year 1 $ $ $ 2 3 4 5 b. Double-declining-balance method Accumulated Depreciation, Year Depreciation Expense End of Year Book Value, End of Year 1 $ $ $ 2 3 4 5 2. On March 4, journalize the entry to record the sale assuming the manager chose the double-declining-balance method. Refer to the Chart of Accounts for exact wording of account titles. PAGE 1 JOURNALACCOUNTING EQUATION DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY 3. On March 4, journalize the entry to record the sale in (2), assuming that the equipment was sold for $103,386 instead of $149,286. Refer to the Chart of Accounts for exact wording of account titles. PAGE 1 JOURNALACCOUNTING EQUATION DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

Please answer in the format I have provided above.

Thank you so much
Answer #1
1(a) STRAIGHT LINE METHOD
Annual Depreciation =(Cost-Residual Value)/Useful Life
Annual Depreciation $160,744 (905600-101880)/5
Depreciation in 5th year , following mid month convention,
Considering 11.5 months of depreciation $154,046 (1607544*(11.5/12)
A B C=905600-B
Year Depreciation Acc. Depreciation Book Value at end of year
1 $160,744 $160,744 $744,856
2 $160,744 $321,488 $584,112
3 $160,744 $482,232 $423,368
4 $160,744 $642,976 $262,624
5 $154,046 $797,022 $108,578
1(b) DOUBLE DECLINING BALANCE METHOD
Annual Depreciation Rate =2*(1/5)= 40%
A B C=A*B D=A-C
Beginning Depreciation Annual Ending Accumulated
Year Book Value Rate Depreciation Book Value Depreciation
1 $905,600 40% $362,240 $543,360 $362,240
2 $543,360 40% $217,344 $326,016 $579,584
3 $326,016 40% $130,406 $195,610 $709,990
4 $195,610 40% $78,244 $117,366 $788,234
5 $117,366 $15,486 $101,880 $803,720
Depreciation will stop when book value equals Estimated Salvage Value
Year 5 depreciation = $15,486
YEARWISE DEPRECIATION
Year Depreciation
1 $362,240
2 $217,344
3 $130,406
4 $78,244
5 $15,486
2 Sales Price $149,286
(Assuming Cash sales)
JOURNAL ENTRY
Date Account Title Debit Credit
March,04 Cash $149,286
Accumulated Depreciation $803,720
Gain on Disposal $47,406
Equipment $905,600
3 Sales Price $103,386
(Assuming Cash sales)
JOURNAL ENTRY
Date Account Title Debit Credit
March,04 Cash $103,386
Accumulated Depreciation $803,720
Gain on Disposal $1,506
Equipment $905,600